OpenStudy (anonymous):

How to solve this polynomial? 3x^3(4x^2)

OpenStudy (anonymous):

\[3x^3(4x^2)= 3*4(x^{3+2})=12x^5\]

OpenStudy (anonymous):

Thankyou so much !

OpenStudy (anonymous):

ur welcme ;)